OpenStudy (toxicsugar22):

According to the following reaction, how many grams of water are necessary to form 0.478 moles oxygen gas? 2 H2O (l) = 2 H2 (g) + O2 (g)

OpenStudy (anonymous):

17.23 grams of H2O \[.478 mol O2\times \frac{ 2 mol H2O }{ 1 mol O2 } = 0.956 mol H2O \times \frac{ 18.02 g H2O }{ 1 mol H2O} = 17.23 g H2O\]

OpenStudy (toxicsugar22):

thanks